# PICASSO: A Feed-Forward Framework for Parametric Inference of CAD Sketches via Rendering Self-Supervision

### v1

@@ 5.2 Few-shot and Zero-shot CAD Sketch Parameterization
Few-shot Evaluation: For the few-shot setting, we first pre-train PICASSO with rendering self-supervision (defined in Eq.(1)). The learned CAD sketch parameterization model is subsequently fine-tuned on smaller, curated sets of parameterized sketches (as discussed in Section. 4.4). Note that two separate models are pre-trained and finetuned for hand-drawn and precise sketch images. In Figure 13, the pre-trained PICASSO (w/ pt.) is compared to its from-scratch counterpart (w/o pt.). Overall, the pre-training outperforms learning from scratch across different sizes of fine-tuning datasets both with precise and hand-drawn images. A comparison with Vitruvion [37] and ResNet34 baseline on a 16k-shot setting is reported in Table 1. We observe that Vitruvion completely underperforms when trained with only a small number of annotated samples. Furthermore, we demonstrate in Figure 5 (right) that PICASSO is able to to parameterize challenging hand-drawn sketches even when fine-tuned with a set of only $2$k annotated samples. To assess the generalization capabilities of all methods, a cross-dataset evaluation is considered, where few-shot models trained on Sketchgraphs are tested on the CAD as a Language dataset [14]. Table 2 shows that PICASSO consistently outperforms Vitruvion [37] and Resnet34 baseline.
Figure 8: Zero-shot setting. Qualitative results of PICASSO learned CAD sketch parameterization from hand-drawn sketches.
Zero-shot Evaluation: By leveraging rendering self-supervision, PICASSO can estimate the parameters of sketches directly without requiring parametric supervision. We evaluate the performance of PICASSO on the challenging zero-shot CAD sketch parameterization scenario in Table 3. Comparison is performed w.r.t. to PMN [4], a recent self-supervised method for vectorized sketch abstraction. PMN operates on strokes (not images) that we form by sampling points on primitives. To adapt PMN for CAD sketch parameterization, we parameterize the predicted drawing primitive (stroke and type) via least-squares fitting. The network is trained on the SketchGraphs dataset [36]. We observe that PICASSO surpasses PMN performance in terms of image-based metrics. Note that we do not report Acc and ParamMSE on the zero-shot setting as any given CAD sketch can be constructed by arbitrarily many parameterizations and a self-supervised method cannot exactly match the parameters employed by the designer (by merging co-linear lines, flipping start-end points, etc.). Figure 8 shows some qualitative results of PICASSO on the zero-shot setting. Note that despite the large complexity of the task, PICASSO still recovers plausible sketch parameterization.
Method ImgMSE CD
PMN [4] 0.233 3.243
PICASSO 0.184 1.880
Table 3: Zero-shot evaluation on hand-drawn images. Comparison is performed w.r.t. the method of [4] that is adapted for CAD sketch parameterization and trained on the SketchGraphs dataset.
@@ 5.3 Ablation Study
SPN with Parametric Supervision: We evaluate SPN of PICASSO for CAD sketch parameterization of precise and hand-drawn sketch images. For this evaluation, SPN is trained with parametric supervision as described in Section 4.2. Results are reported in Table 6. The proposed feed-forward SPN surpasses both the state-of-the-art autoregressive Vitruvion [37] and the non-autoregressive Resnet34 baseline by a large margin. Interestingly, while overall, performance decreases for all methods when trained on hand-drawn sketches, Vitruvion suffers a much larger drop compared to our proposed approach. It is also important to mention that SPN is approximately 10×10$\times$10 × faster than Vitruvion thanks to its feed-forward nature.
@@ 6 Conclusions and Future Works
In this paper, PICASSO is introduced, a novel framework for CAD sketch parameterization. It includes a feed-forward Sketch Parameterization Network (SPN) and a Sketch Rendering Network (SRN). SRN is a neural differentiable renderer that enables the use of rendering self-supervision for large-scale CAD sketch parameterization without requiring parametric annotations. This in turn accounts for real-world scenarios where large annotated hand-drawn CAD sketch datasets are not available. Extensive experiments on few- and zero-shot settings validates the effectiveness of the proposed framework. Due to data unavailability, PICASSO was evaluated on four types of primitives (lines, points, arcs, and circles). In the supplementary, preliminary experiments are conducted with synthetic b-splines suggesting the applicability of PICASSO to free-form curves. This investigation in the context of CAD is left for future. Future work will also address joint CAD sketch and constraint parameterization.
@@ 7 Acknowledgements
The present work is supported by the National Research Fund (FNR), Luxembourg under the BRIDGES2021/IS/16849599/FREE-3D project and Artec3D.
@@ 12 Semi-Supervised CAD Sketch Parameterization via Rendering Supervision
Method Acc ParamMSE ImgMSE CD
PICASSO (w/o pt.) 0.595 451 0.156 2.789
PICASSO (w/o pt. + semi-supervised) 0.608 432 0.145 1.833
Table 10: Semi-supervised learning results for PICASSO.
13 Extension to Other Primitives:
Method CD
SPN 1.07
SPN+DiffVG 3.60
SPN+SRN 0.48
Figure 18: PICASSO on synthetic CAD sketch including B-spline.

### v2

@@ 5.1 Few-shot and Zero-shot CAD Sketch Parameterization
Few-shot Evaluation: For the few-shot setting, we first pre-train PICASSO with rendering self-supervision (defined in Eq.(1)). The learned CAD sketch parameterization model is subsequently fine-tuned on smaller, curated sets of parameterized sketches (as discussed in Section. 4.4). Note that two separate models are pre-trained and finetuned for hand-drawn and precise sketch images. In Figure 15, the pre-trained PICASSO (w/ pt.) is compared to its from-scratch counterpart (w/o pt.). Overall, the pre-training outperforms learning from scratch across different sizes of fine-tuning datasets both with precise and hand-drawn images. A comparison with Vitruvion [38], PpaCAD [40], and ResNet34 baseline on a 16k-shot setting is reported in Table 1. We observe that both Vitruvion and PpaCAD completely underperform when trained with only a small number of annotated samples. Furthermore, we demonstrate in Figure 5 (right) that PICASSO is able to to parameterize challenging hand-drawn sketches even when fine-tuned with a set of only $2$k annotated samples. To assess the generalization capabilities of all methods, a cross-dataset evaluation is considered, where few-shot models trained on Sketchgraphs are tested on the CAD as a Language dataset [14]. Table 2 shows that PICASSO consistently outperforms Vitruvion [38], PpaCAD [40], and the Resnet34 baseline.
Zero-shot Evaluation: By leveraging rendering self-supervision, PICASSO can estimate the parameters of sketches directly without requiring parametric supervision. We evaluate the performance of PICASSO on the challenging zero-shot CAD sketch parameterization scenario in Figure 8 (left). Comparison is performed w.r.t. to PMN [4], a recent self-supervised method for vectorized sketch abstraction. PMN operates on strokes (not images) that we form by sampling points on primitives. To adapt PMN for CAD sketch parameterization, we parameterize the predicted drawing primitive (stroke and type) via least-squares fitting. The network is trained on the SketchGraphs dataset [37]. We observe that PICASSO surpasses PMN performance in terms of image-based metrics. Note that we do not report Acc and ParamMSE on the zero-shot setting as any given CAD sketch can be constructed by arbitrarily many parameterizations and a self-supervised method cannot exactly match the parameters employed by the designer (by merging co-linear lines, flipping start-end points, etc.). Figure 8 (right) shows some qualitative results of PICASSO on the zero-shot setting. Note that despite the large complexity of the task, PICASSO still recovers plausible sketch parameterization.
Method ImgMSE CD PMN [4] 0.233 3.243 PICASSO 0.184 1.880 Refer to caption
Figure 8: Zero-shot evaluation on hand-drawn images. (left) Comparison w.r.t. the method of [4] adapted for CAD sketch parameterization and trained on the SketchGraphs dataset. (right) Qualitative results for PICASSO learned in a zero-shot setting.
@@ 5.2 Ablation Study
SPN with Parametric Supervision: We evaluate SPN of PICASSO for CAD sketch parameterization of precise and hand-drawn sketch images. For this evaluation, SPN is trained with parametric supervision on the complete SketchGraphs dataset. Results are reported in Table 5. SPN surpasses both the autoregressive Vitruvion [38] and the non-autoregressive Resnet34 baseline by a large margin. It is also important to mention that SPN is approximately 10×10$\times$10 × faster than Vitruvion thanks to its feed-forward nature.
@@ 6 Conclusions and Future Works
In this paper, PICASSO is introduced, a novel framework for CAD sketch parameterization. It includes a feed-forward Sketch Parameterization Network (SPN) and a Sketch Rendering Network (SRN). SRN is a neural differentiable renderer that enables the use of rendering self-supervision for large-scale CAD sketch parameterization without requiring parametric annotations. This in turn accounts for real-world scenarios where large annotated hand-drawn CAD sketch datasets are not available. Experiments on few- and zero-shot settings validate the effectiveness of the proposed framework. Due to data unavailability, PICASSO was evaluated on four types of primitives (lines, points, arcs, and circles). In the supplementary, preliminary experiments are conducted with synthetic b-splines suggesting the applicability of PICASSO to free-form curves. This investigation in the context of CAD is left for the future.
@@ 7 Acknowledgements
This work is supported by the National Research Fund (FNR), Luxembourg, BRIDGES2021/IS/16849599/FREE-3D project and by Artec3D.
@@ 10 Implementation Details for Comparative Analysis
PpaCAD: We implemented the concurrent method PpaCAD [40] as no public code is available. Image encoding uses a 2-layer patch MLP followed by a transformer. Separate losses are computed for each primitive type, parameter, and construction flag. Further details are in  [40].
@@ 11 Semi-Supervised CAD Sketch Parameterization via Rendering Supervision
Method Acc ParamMSE ImgMSE CD
PICASSO (w/o pt.) 0.595 451 0.156 2.789
PICASSO (w/o pt. + semi-supervised) 0.608 432 0.145 1.833
Table 9: Semi-supervised learning results for PICASSO.
12 Sensitivity to Rendering Quality
For PICASSO, rendering self-supervision is facilitated by neural differentiable rendering via SRN. We conduct an ablation study to explore how variations in SRN’s rendering performance influence the effectiveness of self-supervised pretraining. To that end, we vary SRN rendering quality by training SRN for different number of epochs. Specifically, PICASSO is pretrained via rendering self-supervision using different SRN renderers trained for 5, 10, 30 and 40 epochs. The results are presented in Figure 13. We find that SRN rendering performance (measured in terms of chamfer distance) stabilizes after a few training epochs (orange line). However, the zero-shot performance of SPN is notably stronger when using a fully trained SRN (blue line). As the neural rendering improves, fewer artifacts are introduced and SRN can more accurately replicates the input sketch parameters.
Figure 13: (orange) Rendering performance (in terms of chamfer distance) of SRN trained for different number of epochs. (blue) Zero-shot performance of PICASSO-SPN when pretrained by the aforementioned SRN renderers.
13 Extension to Other Primitives
Method CD
SPN 1.07
SPN+DiffVG 3.60
SPN+SRN 0.48
Figure 14: PICASSO on synthetic CAD sketch including B-spline.
@@ 14.3 Test-time Optimization with SRN
15 Permutation Invariance of SRN w.r.t. Primitive Order
Since SRN processes input primitive tokens as a set, it should demonstrate permutation invariance with respect to their order. Although this invariance is not explicitly designed into the architecture, it naturally emerges from the synthetic training process, where primitives can appear in any sequence position. Figure 20 qualitatively illustrates the model’s robustness to input order, with renderings of CAD sketches showing only subtle differences despite permuted primitive sequences.
Figure 20: SRN renderings of the parametric CAD sketches where the primitives are randomly permuted.
16 Failure Case Analysis
We conduct a qualitative analysis of common failure cases identified for PICASSO finetuned with 16k samples. Figure 21 highlights several of these cases, where the model produces suboptimal parameterizations. Notable issues include difficulty handling closely positioned primitives, inaccurate coordinate predictions, missing primitives, failure to capture fine details, overly large arc predictions or combinations of the aforementioned cases.
Figure 21: Common failure cases of PICASSO finetuned with parametric supervision on 16k samples.
17 PICASSO for Feature-based 3D CAD Modeling
2D CAD sketches are an essential component of feature-based CAD modeling. In Figure 22, we demonstrate how PICASSO enables the design of 3D CAD models by parameterizing hand-drawn sketches. These parameterized sketches are then uploaded to Onshape [2], where simple extrusions and revolutions are applied. This results in 3D solids that are combined to form the 3D models depicted in the figure.
Figure 22: 3D CAD modeling from hand-drawn sketches with PICASSO.